OpenStudy (anonymous):

can anyone help me out with a few problems please here is one (solve for x: 5x – 4 – x = -3x – 6 – x

OpenStudy (anonymous):

another is -6x – 2 = 40

OpenStudy (anonymous):

5x-4-x=-3x-6-x 4x-4=-4x-6 4x+4x=-2 8x=-2 x=-1/4

OpenStudy (anonymous):

-6x-2=40 6x=-2-40 6x=-42 x=-7

OpenStudy (anonymous):

6( x – 4 ) = 3( 3x + 5)

OpenStudy (anonymous):

6(x-4)=3(3x+5) 6x-24=9x+15 9x-6x=-24-15 3x=-39 x=-13

OpenStudy (anonymous):

Do you know how to do Graph each equation:

OpenStudy (anonymous):

sorry I ment graph equations

OpenStudy (anonymous):

I only have 3 problems to graph and I have no clue how to do them

OpenStudy (anonymous):

Graphing them is easy all you need to do is find the point on either the y or x axis z. Okay like the x= -13 you find negative on the x axis then you make a line straight across its that easy

OpenStudy (anonymous):



OpenStudy (anonymous):

sorry thats a -13